OpenStudy (anonymous):

Simplify the expression using distributive property 2t + 5 (1-4t) =

OpenStudy (hunus):

If a(b + c) = ab + ac What does 5(1 - 4t) equal?

OpenStudy (hunus):

5(1 - 4t) = 5(1 + (-4)t)